A 9.168 M aqueous solution of H2SO4 has a density of 1.4987 g/ml at 20

 C. How many grams of H2SO4 are in 50.0 g of this solution? 


1
Expert's answer
2022-03-30T02:22:04-0400

Solution:

Molarity = Moles of solute / Liters of solution


Calculate the liters of solution:

Density = Mass / Volume

Therefore,

Volume = Mass / Density = (50.0 g) / (1.4987 g/ml) = 33.36 mL

Liters of solution = (33.36 mL) × (1L / 1000 mL) = 0.03336 L


Calculate the moles of solute (H2SO4):

Moles of H2SO4 = Molarity × Liters of solution = (9.168 M) × (0.03336 L) = 0.3058 mol


Calculate the mass of solute (H2SO4):

The molar mass of H2SO4 is 98.079 g/mol

Therefore,

Mass of H2SO4 = (0.3058 mol H2SO4) × (98.079 g H2SO4 / 1 mol H2SO4) = 29.99 g H2SO4

Mass of H2SO4 = 30.0 g


Answer: 30.0 grams of H2SO4 are in this solution